A firm evaluates all of its projects by applying the NPV decision rule. A project under consideration has the following cash flows:

Year Cash Flows
0 -3,002
1 1,036
2 2,229
3 1,649

What is the NPV for the project if the required return is 25 percent?
Answer #1

NPV is the difference between the present value of cash inflows and the present value of cash outflows over a period of time.

CF CF CF3 *0+ (1+r)1*(1 + r)2 + (1+r)3

NPV = -3002 + 1036 2229 1649 (1 +0.25)1 *(1 +0.25)2 + (1 +0.25)3

NPV = -3002 + 828.80 + 1426.56 +844.29 = 97.65